I need to find the starting number to my question:
X times 7, +20, then divided by 30 to equal 10
thankyou

Answer by rothauserc(4718)

((x * 7) +20) / 30 = 10
(7x + 20) / 30 = 10
multiply both sides of = by 30
7x + 20 = 300
7x = 280
x - 40